Please help me find the slope and y intercept of -5x + 7y = 2
Thank you

Answer by Septimus314(7)

set it up in the equation of a line which is y=mx + b
m = slope, b = y intercept
7y = 5x + 2, divide both sides by 7
y = (5/7)x + 2/7
m =5/7
b= 2/7